OpenStudy (azionne):

4(3 - 2x) = 15 solve and will medal

pooja195 (pooja195):

1st start by distributing 

OpenStudy (azionne):

ok

pooja195 (pooja195):

Wat is 4 times 3 ?

OpenStudy (azionne):

12

pooja195 (pooja195):

4 times -2 = ?

OpenStudy (azionne):

-8

pooja195 (pooja195):

good so \[\huge~\rm~12-8x=15\] next step subtract 12 from both sides 15-12=?

OpenStudy (azionne):

3

OpenStudy (azionne):

i got 2.6

pooja195 (pooja195):

Thats inccorect

OpenStudy (azionne):

O.o

pooja195 (pooja195):

\[\huge~\rm~8x/-8=3/-8\] We need to divide both sides by -8 8/-8=?

OpenStudy (azionne):

0

pooja195 (pooja195):

1... so you get x=3/-8

pooja195 (pooja195):

BUT

pooja195 (pooja195):

When you simplfy it you get -3/8 so x=-3/8

OpenStudy (azionne):

thx

pooja195 (pooja195):

np